OpenStudy (twistrose):

What is the solution to the equation? 2(x – 7) = 10x + 18

OpenStudy (twistrose):

please help!

OpenStudy (jdoe0001):

what did "you" get anyway?

OpenStudy (twistrose):

i have a hard time with this problem. I was hoping someone could show me?

OpenStudy (anonymous):

2x - 14 = 10x + 18 2x - 10x = 14 + 18 -8x = 32 x = 32/-8 x = -4

OpenStudy (twistrose):

thanks

OpenStudy (anonymous):

for sure 100%

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

can i get a medal

OpenStudy (mathmale):

@mikey356333: Kindly do not ask for medals.

OpenStudy (anonymous):

ok plz

OpenStudy (anonymous):

i just got reported by some one